### v1

# Rejoining fragmented ancient bamboo slips with physics-driven deep learning

Abstract
Bamboo slips are a crucial medium for recording ancient civilizations in East Asia, and offers invaluable archaeological insights for reconstructing the Silk Road, studying material culture exchanges, and global history. However, many excavated bamboo slips have been fragmented into thousands of irregular pieces, making their rejoining a vital yet challenging step for understanding their content. Here we introduce WisePanda, a physics-driven deep learning framework designed to rejoin fragmented bamboo slips. Based on the physics of fracture and material deterioration, WisePanda automatically generates synthetic training data that captures the physical properties of bamboo fragmentations. This approach enables the training of a matching network without requiring manually paired samples, providing ranked suggestions to facilitate the rejoining process. Compared to the leading curve matching method, WisePanda increases Top-50 matching accuracy from 36% to 52%. Archaeologists using WisePanda have experienced substantial efficiency improvements (approximately 20 times faster) when rejoining fragmented bamboo slips. This research demonstrates that incorporating physical principles into deep learning models can significantly enhance their performance, transforming how archaeologists restore and study fragmented artifacts. WisePanda provides a new paradigm for addressing data scarcity in ancient artifact restoration through physics-driven machine learning.
Main
Figure 1: The challenge of bamboo slip rejoining in archaeological research. The excavated bamboo slip fragments from the Qin dynasty (221–206 BCE) arranged in a spiral pattern (left) illustrate the overwhelming scale of the rejoining challenge faced by archaeologists—with thousands of fragments at single excavation sites, each potentially matching with any other piece. A magnified view of a representative fragment (upper right) reveals the characteristic irregular breakage pattern along the fracture edge, while an example of a reconstructed complete slip (lower right) demonstrates how proper rejoining establishes textual content continuity, which is essential for the interpretation of ancient texts. The complex morphology of the fractures, coupled with deterioration from millennia underground, creates a matching problem of extraordinary complexity that has traditionally required intensive manual effort, with experts sometimes spending weeks to successfully match a single pair of fragments.
Bamboo slips, serving as a fundamental medium for documenting ancient East Asian civilizations, contain invaluable historical records spanning philosophy, law, and social life period [1, 2, 3, 4, 5]. Their durability has enabled these artifacts to survive millennia underground while retaining legible content [6], offering scholars unprecedented insights into historical societies. However, the excavation of these delicate artifacts presents a critical challenge [7] - many bamboo slips have been fragmented into thousands of pieces (Figure 1), significantly complicating efforts to reconstruct and interpret their content. This fragmentation creates a fundamental obstacle in accessing the wealth of historical knowledge contained in these artifacts.
The rejoining of fragmented bamboo slips represents one of the most challenging problems in archaeological preservation and cultural heritage studies [8]. The difficulty stems from multiple factors: First, the enormous number of potential fragment combinations makes manual matching extremely time-consuming - for instance, The Qin bamboo slips from the Shuihudi site comprise fragments numbering in the tens of thousands, with each piece potentially matching any of the others [9]. Second, environmental factors like moisture and pressure have caused extensive physical deterioration [10, 11], distorting the original shapes and surfaces. Third, the unique fiber structure of bamboo creates complex fracture patterns that traditional curve-matching approaches struggle to analyze effectively. These challenges have made the restoration process highly labor-intensive, with experts sometimes requiring weeks to successfully match a single pair of fragments [12].
Deep learning for ancient artifact repair
Here we overcome the challenges of fragmented bamboo slip rejoining through physics-driven deep learning. Our approach integrates principles from the physics of fracture with deep neural networks, enabling automated matching of fragmented pieces without relying on manually labeled training data. Deep learning has been successfully used in restoring and attributing ancient texts [13] empowered by large labeled dataset. While recent years have also seen significant progress in fragmented cultural relic restoration [14] with labeled data - from oracle bone reconstruction [15, 16, 15, 17] to manuscript fragment matching [6] - most existing approaches rely on traditional computer vision techniques and struggle with complex degradation patterns [18]. Some works have attempted to use generative models for cultural heritage restoration [17]. While achieving promising results, these methods are constrained by the scarcity of training data and often fail to capture the intricate physical properties of ancient materials [19, 20].
To address these fundamental challenges, we propose WisePanda, a deep learning framework that leverages physical principles to overcome the data scarcity problem inherent in fragmented bamboo slip rejoining tasks. Training such a system presents a unique paradox: while manual fragment rejoining is prohibitively time-consuming - the very problem we aim to solve - this same process would traditionally be required to generate training data for the model. We resolve this dilemma by resorting to the physics of fracture [21, 22]. By modeling the physical properties of bamboo and the process that govern its degradation [23, 10], we generate extensive synthetic training data that captures the essential characteristics of real paired fragment slips (Figure 2). This physics-driven approach enables us to produce large-scale, realistic training data without requiring manual matching efforts, while ensuring the model learns meaningful patterns based on actual material properties rather than superficial features.
Figure 2: Physics-driven modeling of bamboo slip fracture and deterioration. a, The breakage process of bamboo slips showing how fracture propagates across the bamboo’s fiber structure, with the resulting irregular curve (black line) and the corresponding stress field distribution (blue gradient). b, Detailed stress field model illustrating the mathematical relationship between the fracture angles ($\theta_{i-1}$, $\theta_{i}$) and the stress propagation in the $x$-$y$ coordinate system, where the blue dotted circles represent the stress field emanating from the fracture endpoint, showing how stress radiates outward and concentrates at fiber boundaries. c, Time-sequential simulation of the corrosion process, demonstrating how the original fracture pattern (left) changes through environmental exposure (middle) to produce the final deteriorated edge morphology (right), with protruding areas experiencing accelerated degradation. d, Collection of real bamboo slip fracture curves extracted from archaeological samples, exhibiting diverse breakage patterns that serve as reference for model validation. e, Synthetically generated fracture curves produced by our physics-driven model, displaying morphological characteristics highly similar to the real samples in panel d. This approach enables the generation of extensive paired training data that captures both the physical properties of bamboo fragmentation and the effects of long-term degradation, effectively addressing the data scarcity challenge inherent in ancient artifact restoration.
Data synthesis with physics of fracture
The generation of training data for bamboo slip rejoining requires a comprehensive understanding of both fracture formation and degradation processes [24]. Our physical model builds on two key observations: First, bamboo slips exhibit distinct “transverse” and “longitudinal” fracture modes due to their unique fiber structure [24], as illustrated in Extended Figure E1, with transverse fractures accounting for approximately 70% of cases. Second, these fractures undergo complex corrosion processes over time, significantly altering their original patterns [25, 10, 2]. Building on these insights, we develop a systematic approach to generate synthetic fragment pairs for transverse fractures, which represent the majority of rejoining challenges.
The physics of fracture in bamboo slips is governed by their distinctive vertical fiber arrangement [26, 27, 28, 29], as detailed in Extended Figure E2. When stress extends horizontally across the slip, it creates characteristic fracture patterns as it propagates through consecutive fiber bundles [30, 27, 29]. We model this process through a probabilistic framework where the fracture angle of each fiber bundle is influenced by the stress field generated at the fracture endpoint of the preceding fiber bundle. This physical relationship can be expressed through a probability density function that determines likely fracture paths based on stress distribution and geometric configurations. The corrosion process is then simulated by analyzing geometric exposure - protruding areas and exposed surfaces are more susceptible to degradation through moisture absorption and microorganism activity [31]. To elegantly calculate the exposure area of each fiber, we ingeniously employ ReLU functions that compare the height differences between adjacent fibers, effectively capturing only the protruding portions that are more vulnerable to environmental deterioration. This computational approach allows us to precisely model the progressive and differential deterioration patterns observed in archaeological samples.
To ensure the generated data accurately reflects real bamboo slip characteristics, we optimize model parameters using a genetic algorithm [32, 33]. The algorithm evaluates parameter sets by comparing the distribution of generated fragments against a reference set of 200 real fragment edges through dimensionality reduction and Silhouette analysis [34]. Our analysis shows high similarity between the synthetic and real data distributions. The resulting pipeline enables us to generate extensive training data that captures both the physical properties of fracture formation and the effects of long-term degradation, effectively addressing the data scarcity challenge in rejoining fragmented bamboo slips.
WisePanda is a ranking system
Figure 3: Framework of WisePanda integrating physics-driven deep learning for bamboo slip rejoining. The workflow diagram illustrates the complete pipeline of our system. The process begins with input bamboo slip fragments (top row), where the target fragment is highlighted in pink. In the physical model stage, we implement two key components: a physics-driven fracture engine that generates initial breakage patterns, and a corrosion simulation modifier that replicates long-term degradation effects. The dimensionality reduction module (center) enables comparison between real and synthetic fragments through statistical visualization techniques, facilitating parameter optimization to ensure realistic fracture simulation. The optimized physical model feeds into the training pair synthesizer, which generates matching fragment pairs for training the matching network. This matching network learns to effectively represent and compare fragment features through a multi-layer architecture with specialized embedding capabilities. During inference (red arrows), the system compares the target fragment against all candidates, producing a ranked list of potential matches (bottom row). The ranking transforms thousands of possible matches into a manageable set, with the system highlighting the most probable candidates (position 5 shown in pink represents a correct match). This approach enables archaeologists to focus their verification efforts on the most promising candidates, significantly accelerating the traditionally time-consuming rejoining process.
WisePanda’s architecture features an interpretable pipeline for fragmented bamboo slip rejoining task (Figure 3), using synthesized data with physics of fracture for model training and providing top-k candidates to assist archaeologists. At its core lies a TripletNet-based deep learning network [35], designed to learn effective feature representation that can distinguish between matching and non-matching fragments. The network processes each input as a 64-dimensional vector that captures the detailed features of the fracture edge curves [36, 37, 38]. By leveraging a triplet loss function [39, 40, 41], the network learns to minimize the distance between matching pairs while maximizing the distance between non-matching ones.
In practical deployment, WisePanda operates by predicting matching probabilities between fragments and offers archaeologists a ranked list of the top potential matches for each fragment. This significantly reduces the search space from thousands of candidates to a manageable set [42, 43, 44], thereby streamlining the rejoining process. The network’s output is a value between 0 and 1, indicating the model’s confidence in a match between the input fragments, where a value closer to 1 signifies a higher likelihood of a correct match. Moreover, this ranking approach effectively mitigates the overwhelming task of manually matching fragments by narrowing down the possibilities to the most probable matches [45].
WisePanda as an archaeologist adjunct
To bridge the gap between computational methods and archaeological research practices, we have developed an intuitive computer-assisted tool that implements WisePanda’s capabilities in a practical workflow (Extended Figure E5). When working with this tool, archaeologists first select a target fragment from their collection. The system then automatically generates a ranked list of potential matching candidates [43, 42, 46], significantly narrowing down the search space from thousands to dozens of fragments. Each suggested match is accompanied by a confidence score derived from WisePanda’s matching network, allowing archaeologists to prioritize their examination of the most promising candidates.
The tool provides a dedicated workspace where archaeologists can visually manipulate and compare fragments through intuitive operations such as dragging, rotating, and fine-tuning alignments. To facilitate detailed examination, the system offers specialized visualization features like layer swapping and edge enhancement, enabling careful inspection of ink traces and fracture patterns All successful matches are saved and documented in a centralized database, creating a growing repository of verified rejoining that contributes to the broader restoration effort. This computer-assisted tool has significantly accelerated the traditional rejoining workflow [47, 44], ensuring archaeologists maintain full control over the final verification of matches while benefiting from AI-powered suggestions.
Experimental evaluation
Datasets. For evaluating WisePanda’s effectiveness, we compiled several datasets of bamboo and wooden slip fragments. Our primary test set (Bamboo236) consists of 118 pairs of matched bamboo slip fragments, carefully verified and provided by archaeological experts who successfully rejoined these fragments through meticulous manual work. These bamboo slip artifacts are part of an ongoing archaeological research project, with the cultural relics yet to be formally published. To simulate more challenging real-world scenarios, we created an extended dataset (Bamboo1350) by introducing 1114 additional interference fragments, expanding the candidate pool to 1,232 total fragments while maintaining the original test pairs. To assess generalization capabilities across different materials, we also compiled wooden slip datasets. The base wooden slip test set (Wood670) contains 335 expert-verified pairs, while its extended version (Wood3833) includes 3163 additional interference fragments. These diverse datasets enable comprehensive evaluation of fragment matching methods under varying conditions and material types.
Evaluation methodology. The evaluation process involved computing similarity scores between each fragment’s fracture edge and all potential matching candidates from the complementary group. Fragments were pre-classified as upper or lower parts, with matching only performed between these distinct groups. We extracted features from both fracture edges of each fragment, ranked the computed similarity scores, and measured the probability of finding the correct match within the Top-k predictions. We conducted comprehensive experiments comparing our approach with several existing methods, including Manual Random Search, curve matching techniques (using Cosine Similarity and Dynamic Time Warping (DTW)) [48], Fast Matching Method (FMM) [49, 50], and Scale Invariant Signature (SIS) [51]. This comparative framework allows us to assess the relative performance of our physics-driven approach against established techniques in the field of fragment matching.
Results and analysis. The performance variations among different methods reveal insights about the nature of the bamboo slip matching problem. In our original test set (Bamboo118), WisePanda demonstrated superior performance across all metrics, achieving a top-1 accuracy of 12.29%, top-5 accuracy of 35.17%, top-10 accuracy of 52.54%, top-20 accuracy of 69.07%, and a top-50 accuracy of 94.07%. Among traditional methods, SIS achieved the second-best performance with a top-1 accuracy of 10.59% and top-50 accuracy of 72.88%, followed by DTW with a top-50 accuracy of 75.42%. FMM and Cosine Similarity performed less effectively, with top-50 accuracies of 65.25% and 17.80% respectively.
To simulate real-world archaeological scenarios where numerous unrelated fragments exist, we expanded our test dataset by introducing additional interference fragments. As shown in Extended Figure E4, when the candidate pool was increased by a factor of approximately 10, WisePanda maintained a top-50 accuracy of 52.54% compared to the original 94.07%, demonstrating robust performance even with significantly increased search space.
Method Top1(%) Top5(%) Top10(%) Top20(%) Top50(%) Top100(%)
Bamboo236 Manual Random Search 0.80 4.22 8.40 17.00 42.30 -
DTW [48] 7.20 22.46 30.93 49.58 75.42 -
FMM [50] 8.90 18.64 26.27 38.14 65.25 -
SIS [52] 10.59 22.03 30.93 47.03 72.88 -
WisePanda 12.29 35.17 52.54 69.07 94.07 -
Bamboo1350 Manual Random Search 0.08 0.44 0.89 1.79 4.48 8.97
DTW 3.81 12.29 14.41 20.34 32.20 37.71
FMM 6.36 11.44 13.14 16.10 25.00 33.47
SIS 7.20 15.25 19.07 24.15 36.02 48.73
WisePanda 5.93 18.22 24.58 37.29 52.54 66.10
Wood WisePanda[Wood670] 5.38 15.37 26.42 38.36 64.18 86.27
WisePanda[Wood3833] 2.69 9.55 13.88 21.64 34.78 48.36
Table 1: Comparison of fragment matching methods across different test scenarios. Results show Top-k matching accuracy (percentage of correctly identified matches within the top k candidates) across bamboo and wooden slip datasets. The original bamboo test set (Bamboo118) contains 118 paired fragments, while the extended version (Bamboo1114) includes 1,114 fragments with additional interference fragments to create a more challenging matching environment. Similarly, for wooden slips, we compare performance on the original test set (Wood335) and an extended version (Wood3163). As demonstrated in Extended Figure E4, WisePanda consistently outperforms traditional curve matching approaches in both material types and across varying complexity levels, demonstrating the effectiveness of our physics-driven approach in handling diverse archaeological fragments.
Interpretability of WisePanda
The fundamental challenge in generating training data for bamboo slip matching lies in the irreversible nature of the physical degradation process [10, 18]. When archaeologists uncover bamboo slip fragments, they observe only the final degraded edges [7]. The original fracture patterns, which are crucial for accurate matching, are permanently altered by centuries of degradation and cannot be directly inferred from the preserved fragments [8, 53]. This irreversibility poses a significant limitation for conventional data-driven approaches [54, 19, 20, 17], which typically require paired samples of ”before-degradation” and ”after-degradation” states for effective training, data that is inherently unavailable in archaeological contexts. Even generative models like GANs suffer from distribution bias when learning only from successfully matched fragments, failing to incorporate the underlying physical principles that govern fracture formation and degradation.
Our physics-driven model addresses this challenge by modeling the forward process of fragmentation and degradation. We first simulate the initial fracture patterns based on the physics of bamboo fiber structures [24, 55, 56, 23], then apply physical degradation processes to these patterns. This sequential modeling captures both the immediate fracture characteristics and their subsequent transformation through environmental exposure. The physical processes governing bamboo deterioration follow specific geometric principles, where protruding areas and exposed surfaces are more susceptible to degradation [10, 18]. By incorporating these physical constraints, WisePanda generates realistic degradation patterns that closely mirror natural deterioration processes, effectively bridging the gap between the unknown original fracture patterns and the observed degraded edges.
These physical principles not only enhance the quality of synthetic training data but also provide interpretability to our WisePanda’s behavior. The combination of authentic fracture patterns and physically-based degradation enables WisePanda to capture fundamental characteristics of bamboo slip fragmentation, generating abundant realistic training data and leading to more robust and generalizable performance. Unlike pure data-driven approaches that may overfit to superficial patterns in limited training samples, our physics-driven model learns meaningful features grounded in the actual physical processes of fracture and degradation, allowing it to generalize effectively to unseen fragments with diverse deterioration patterns.
Rejoin fragmented ancient wooden slips
To examine WisePanda’s generalization capability, we applied our approach to wooden slips - another important medium for ancient text preservation that shares similar characteristics with bamboo slips but exhibits distinct material properties. While wooden slips served similar documentary purposes, their different fiber arrangement and material structure present unique challenges for fragment matching [57, 58]. We evaluated our method on a dataset of 335 pairs of wooden slip fragments, following the same experimental protocol as with bamboo slips.
As shown in Table 1, WisePanda demonstrated promising cross-material generalization with 64.18% top-50 accuracy on wooden slips. When tested with extended interference fragments in both materials, performance predictably decreased to 52.54% for bamboo slips (Bamboo1350) and 34.78% for wooden slips (Wood3833).
This performance difference can be attributed to two key factors. First, the fiber structure of wooden slips differs substantially from bamboo slips, leading to distinct fracture patterns that deviate from our physics-driven model’s assumptions [59, 60, 26]. Second, the degradation process varies significantly due to diverse preservation conditions - factors such as burial duration, soil moisture levels, microbial activity, and regional soil composition all affect how materials deteriorate over time [61, 62, 63]. Even artifacts of the same material excavated from different sites or time periods may exhibit varying degradation patterns, requiring different physical parameters for accurate modeling. These variations are particularly pronounced when comparing wooden and bamboo materials, as their distinct physical properties interact differently with environmental factors, leading to material-specific degradation processes that need to be carefully considered in the modeling approach.
Discussion
In this work, we present WisePanda, the first physics-driven deep learning framework designed for rejoining fragmented ancient bamboo slips. By incorporating principles from the physics of fracture [64, 65, 27] into the generation of synthetic training data, our approach overcomes the fundamental challenge of data scarcity in cultural heritage preservation. The effectiveness of this approach is demonstrated through superior matching accuracy and the development of practical tools that are now assisting archaeologists in their reconstruction work. Our framework’s success in handling both fracture pattern generation and material degradation simulation highlights the advantage of combining physical principles with deep learning, particularly in scenarios where traditional data-driven approaches are constrained by limited training samples.
Beyond its current implementation, this research opens new horizons in bamboo slip reconstruction and broader cultural heritage preservation [66]. For bamboo slips specifically, our framework can be extended to handle more complex scenarios by incorporating additional physical parameters that account for regional variations and different preservation conditions [67, 68]. Furthermore, our approach establishes a new paradigm for fragment matching where training data is scarce. The methodology can be adapted to other archaeological materials such as wooden slips, ceramics, metals, and pottery [69, 70, 71], by modeling their specific physics of fracture and degradation processes. This adaptability stems from the fundamental nature of our approach while artifacts may vary, the underlying physical principles governing their fragmentation and degradation remain consistent and modelable [64, 72].
In conclusion, WisePanda represents not merely a technical advancement in bamboo slip rejoining, but rather a transformative approach to preserving and studying cultural heritage [73]. By combining physical principles with artificial intelligence, our method provides a robust solution when paired training data is limited but physical principles are well understood. This synergy of physics and deep learning opens new possibilities for archaeological fragment matching, potentially revolutionizing how we recover and interpret artifacts from our past [74, 75].
Code availability
The source code for WisePanda framework is publicly available on GitHub at https://github.com/zhujinchi/wisepanda. The repository contains the implementation of our physics-driven deep learning model for bamboo slip fragment matching, including the fracture modeling algorithms, corrosion simulation processes, neural network architecture and user interface. Users can install the required dependencies by following the installation instructions provided in the repository. The codebase is designed to be accessible to researchers, featuring an intuitive graphical user interface for fragment selection, comparison, and verification with integrated AI assistance. Additionally, we have established a project website at https://wisepanda.tech/ that provides demonstration videos, visual examples, and detailed documentation of the fragment rejoining and visualization tools discussed in this paper. This open-source release aims to facilitate further research in digital archaeology and enable the application of our methods to other types of archaeological fragment matching challenges.
Extended Figure
Figure E1: Typical fracture patterns of ancient bamboo slips. (a), The upper row displays bamboo samples with transverse fractures, the most common fracture mode (approximately 70%), characterized by irregular wavy breakage patterns resulting from stress propagating perpendicular to the bamboo fiber direction. These fractures often have limited text preservation at the fracture edges, with the fragmented characters providing insufficient information for reliable matching, thus making the morphological features of these edges the primary basis for rejoining. (b), The lower row exhibits longitudinal fractures, a less common but significant fracture mode, characterized by relatively smooth fracture edges that occur along the bamboo fiber direction. With these smoother fracture patterns, text matching becomes the primary criterion for correct rejoining, as the physical interlocking features are less pronounced. This diversity of fracture patterns and preservation states highlights the intrinsic complexity inherent in ancient bamboo slip rejoining.
Figure E2: Physics-driven model of bamboo fracture propagation. a, Bamboo microstructure showing vertical fiber bundles (red blocks) and a typical fracture path (dotted black line) traversing across fibers. The black arrows indicate shear forces ($\sigma^{\infty}Z^{+}$) inducing Mode III fracture. b, Bamboo fracture modeling with stress field propagation between fibers. Points $P_{1}$, $P_{2}$, and $P_{3}$ represent positions on consecutive fibers $f_{i-1}$, $f_{i}$ and $f_{i+1}$, with angles $\theta_{i-1}$ and $\theta_{i}$ showing fracture directions of fibers $f_{i-1}$and $f_{i}$. Blue dotted circles illustrate the stress field emanating from point $P_{2}$, determining the probability distribution of potential fracture paths.
Figure E3: Data distribution analysis and matching performance. a, t-SNE distributions of fragment features showing the relationship between synthetic data (blue) and real bamboo fragments (green) across four parameter configurations. Progressive convergence indicates effective parameter optimization in our model. b, Fracture matching quality matrices for 118 real fragment pairs. Deeper blue indicates higher similarity, with diagonal elements representing correct matches. The lower-right matrix shows optimal discriminative performance with clear contrast between prominent diagonal (blue) and off-diagonal elements (white), corresponding to the most effective parameter setting (lower-left t-SNE plot). This confirms that generating more realistic synthetic data directly improves fragment matching accuracy.
Figure E4: Robustness analysis of fragment matching algorithms with increasing interference fragments. Performance comparison across six accuracy levels (Top-1 to Top-100) as candidate pool expands from 0 to 1114 fragments. WisePanda (orange bars) maintains relatively stable performance despite increasing interference, while Manual Random Search (dotted purple line) shows rapid accuracy decline. Other methods (DTW, SIS, FMM, and Cosine Similarity) also exhibit significant performance drops. At maximum interference, WisePanda maintains 50-65% accuracy at Top-50 and Top-100 levels, while competing methods fall below 35%, demonstrating physics-driven approach’s advantage in real archaeological scenarios with thousands of potential matches.
Figure E5: Practical application of WisePanda. a, WisePanda software interface showing two successful matching cases: target bamboo slip fragments (left) with their corresponding matching results (right), demonstrating the system’s verification capabilities. b, Efficiency comparison between WisePanda-assisted and traditional manual matching across five test cases. The table records target and result fragments, system recommendation rank, and time measurements for both methods. WisePanda reduces average matching time from 266.54 to 11.52 seconds—approximately a 20-fold efficiency improvement. In three of five cases, the correct match received top ranking (rank 1), confirming the algorithm’s accuracy. These results validate the physics-driven approach’s practical value for large-scale bamboo slip fragment rejoining.